I HOPE YOU CAN HELP
USE THE SUBSTITUTION METHOD TO SOLVE THE LINEAR SYSTEM
1. X + 2 = -5
4x -3y = 2

2. 3x - 2y =4
x + 3y = 5

3. 3x + y = -2
x + 3y = 2

Answer by tutorcecilia(2152)

Solve for "x" in either problem.
X + 2 = -5
X +2 – 2 = -5 -2
X = -7
.
Plug x=7 back into either equation.
4(-7) -3y = 2
Solve for y
-28 – 3y =2
-28 + 28 -3y = 2+ 28
-3y =30
-3y/-3 = 30/-3
Y = -10
.
So we think that both equations have the values of x = -7
and y = -10.
.
Plug the values back into the original equations to verify:
X + 2 = -5
(-7) + 2 = -5
-7 + 2 = -5
-5 = -5 this checks out
4x -3y = 2
4(-7) – 3(-10) = 2
-28 + 30 = 2
2 = 2 This answer also checks out, so the x and y values are correct